OpenStudy (anonymous):

What is the limit of ( x^2 - 1/x) as x approaches 0 from negative infinite ?

OpenStudy (anonymous):

\[\lim_{x\to 0^{-}}\frac{x^2-1}{x}\]

OpenStudy (anonymous):

yeah that

OpenStudy (anonymous):

\[\infty\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[x \to 0^{-} \implies \frac{1}{x} \to -\infty\]

OpenStudy (anonymous):

x is negative. hence, less than 1 making \((x^2 -1 )\) negative

OpenStudy (anonymous):

thanx